A balloon is filled with helium and let go. It rises into the air accelerating at a rate of 0.151 m s−2.

If the density of air is 1.22 kg m−3 then what is the average density of the balloon? (to 3 s.f and in kg m−3)

(Note: use g = 9.81 m s−2 and ignore any "drag" forces on the balloon such as air resistance.)
1
Expert's answer
2020-04-14T09:39:44-0400
"ma=\rho_a gV-mg"

"\rho_b Va=\rho_a gV-\rho_b gV"

"\rho_b (a+g)=\rho_a g"

"\rho_b (0.151+9.81)=1.22 (9.81)"

"\rho_b=1.20\frac{kg}{m^3}"